OpenStudy (mandii):

What are the characteristics of a radical equation? (1 point) How is solving radical equations similar to solving linear equations? (1 point) Why is it important to check the solutions to a radical equation? (1 point) Create your own radical equation. Describe in complete sentences and demonstrate the process in finding its solution(s). (2 points)

OpenStudy (anonymous):

They are actually to the one half power. You can take a factor in the radical and sqrt it and put in on the outside... Ex. sqrt(28) = sqrt(4 * 7) = sqrt(22 * 7) = 2sqrt(7) sqrt(28) = 2 * sqrt(7) Read more: http://wiki.answers.com/Q/What_are_the_characteristics_of_a_radical_equation#ixzz1XetIGbus